OpenStudy (anonymous):

-3x - 8 + 6x = 7 solve for x someone help plz

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle -3x - 8 + 6x = 7 }\) first add like terms on the left side, I will color them in blue for you. \(\large\color{black}{ \displaystyle \color{blue}{-3x} - 8 \color{blue}{+ 6x} = 7 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{blue}{ -3x + 6x } = ?? }\) (can you tell me ?)

OpenStudy (anonymous):

-3x+6x would be 3x I think

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

what would I do next

OpenStudy (solomonzelman):

now, after you have \(\large\color{black}{ \displaystyle 3x - 8 = 7 }\) add 8 to both sides. \(\large\color{black}{ \displaystyle 3x - 8 \color{red}{+8} = 7 \color{red}{+8} }\) wha will your equation then be ?

OpenStudy (anonymous):

8 and 8 is 16 and 7+8 is 15

OpenStudy (solomonzelman):

you have a negative 8 on the left side, and `-8+8` is equal to what /

OpenStudy (solomonzelman):

?

OpenStudy (anonymous):

it would be -8 cuz a negative and a positive is a negative

OpenStudy (anonymous):

the answers are x= -7/5, x= -1/5, x= -1/3, x=5, and x=45

OpenStudy (solomonzelman):

-8+8= 0, no?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

so, you get 3x-8+8=7+8 3x=15

OpenStudy (solomonzelman):

then divide both sides by 3

OpenStudy (anonymous):

15 divided by 3 is 5

OpenStudy (anonymous):

so the answer would be 5 or nah